a bag contains 2/3 of green marbles if there are 120 marbles that are green how many marbles are in the bag
Answer by shailesh tripathi(2)

if there are 2/3rd of green ,total greens are 120 then total » 2/3 =120/x
» x=120*3/2
»x=180.